HOT-141: Chief Scientist Report
Chief Scientist: T. GREGORY
HOT-141: Chief Scientist Report
CRUISE ID: W0210A
Departed: Nov. 2, 2002 at 0900 (HST)
Returned: Nov. 6, 2002 at 0330 (HST)
Vessel: R/V Wecoma
Operator: Oregon State University
Master of the Vessel: Captain Danny Arnsdorf
Chief Scientist: Thomas Gregory
Marine Technician: Daryl Swensen
1. SCIENTIFIC OBJECTIVES
The objective of this cruise was to continue building a collection of
hydrographic and biogeochemical data at the Hawaii Ocean Time-series
(HOT) stations. Three stations were to be occupied during the cruise,
in the following order:
1) Station 1, referred to as Station Kahe, is located at 21° 20.6'N, 158
° 16.4'W and was to be occupied on Nov. 2 for about 3 hours.
2) Station 2: ALOHA (A Long Term Oligotrophic Habitat Assessment) is
defined as a circle with a 6 nautical mile radius centered at 22° 45'N,
158°W. This is the main HOT station and was to be occupied for 3 days
from Nov. 3 to Nov. 5.
3) Station 6, referred to as Station Kaena, is located off Kaena Point
at 21° 50.8'N, 158° 21.8'W and was to be occupied on Nov. 5 for about 4
A single CTD cast was to be conducted at Station 1 to collect
continuous profiles of various physical and chemical parameters. Water
samples were to be collected at discrete depths for biogeochemical
measurements. PRR and TSRB measurements were also to be made.
Upon arrival at Station ALOHA, net tows were to be conducted followed
by the deployment of a free-drifting sediment trap array. After
deployment, a full-depth CTD cast was to be conducted followed by CTD
casts at strict 3-hour intervals for at least 36 hours for continuous
and discrete data collection followed by another full-depth CTD cast.
The primary production array was to be deployed on Nov. 4 for 12
hours. Plankton net tows were to be conducted near noon and midnight
on Nov. 3 and 4 at Station ALOHA. PRR, TSRB and AC-9/FRRf operations
were to be done around noon Nov. 4 and 5 and a nighttime AC-9/FRRf cast
was to be executed at 0300 on Nov. 5. The drifting sediment trap array
was to be recovered near dawn and the ATE sampler was to be deployed
just before noon on Nov. 5.
Following Station ALOHA operations, the ship was to transit to Station
6. A near-bottom CTD cast (~2500 m) was to be conducted at Station 6
including salinity samples for calibration, after which the ship was to
transit back to Snug Harbor.
The following instruments were to collect data throughout the cruise: a
shipboard ADCP, a thermosalinograph and fluorometer, and an
2. SCIENCE PERSONNEL
Shimi Rii Research Associate UH
Daniel Fitzgerald Research Associate UH
Mark Valenciano Electronics Technician UH
Fernando Santiago-Mandujano Research Associate UH
Thomas Gregory (Chief Scientist) Research Associate UH
Lance Fujieki Computer Specialist UH
Paul Morris Technician UH
Tara Clemente Research Associate UH
Cecilia Sheridan Graduate Student UH
Jennifer Brum Graduate Student UH
Dan Sadler (Watch Leader) Research Associate UH
Anne Gasc Research Associate UH
3. GENERAL SUMMARY
All operations at Stations Kahe and ALOHA were conducted as planned.
The cast at Station Kaena was abbreviated due to a medical emergency.
Thirteen 1000 m and two 4800 m CTD casts were completed at Station
ALOHA. One 1000 m cast at Stations Kahe and one 2500 m cast at Station
Kaena (albeit with fewer bottles tripped) were obtained. Both
free-floating arrays were deployed and recovered without incident. All
optics operations were conducted as planned.
C. Sheridan successfully completed six plankton net tows.
Weather conditions were favorable throughout the cruise.
The ADCP ran without interruption throughout the cruise, as well as the
fluorometer, thermosalinograph and the ship's anemometer.
We arrived back at Snug Harbor on November 6 at around 0330. A
complete off-load took place later that morning.
4. R/V WECOMA, OFFICERS AND CREW, TECHNICAL SUPPORT
The R/V WECOMA and her crew delivered exceptional ship support for our
work. The officers and crew were most helpful and accommodating and are
to be commended for maintaining high standards. They showed enthusiasm
and concern for our work and were very flexible in receiving changes in
our operational schedule. The marine technician provided invaluable
assistance with our operations.
5. DAILY REPORT OF ACTIVITIES (HST)
October 31, 2002; Loading Day
Equipment loaded on this day. The CTD was terminated and tested.
November 2, 2002
The ship departed from Snug harbor at 0900. Fire and abandon ship
drills were conducted at 0930, followed by a short science meeting
during which the cruise schedule was reviewed and safety issues were
We arrived at Station Kahe at 1145 and immediately conducted a weight
cast followed by PRR and TSRB operations and then a 1000 m CTD cast.
The package was back on deck at 1425 and we then began transit to
November 3, 2002
We arrived at Station ALOHA on schedule and immediately performed a net
tow followed by deployment of the sediment trap array. The deep PO
cast started at 0145 and was back on deck at 0542. Following the deep
cast we performed the shallow PO cast, which initiated the 36-hr CTD
cast period. We conducted six 1000 m casts this day.
Net tows were conducted at 0017, 1000, 1302 and 2158.
November 4, 2002
Seven 1000 m CTD casts were conducted this day. The second deep cast
was begun at 2250.
Net tows were performed at 0011 and 0956.
The PRR and TSRB were deployed at 1204. One AC-9/FRRf cast was
conducted at 1243.
The primary production array was deployed at 0633 and recovered at
November 5, 2002
The deep cast was recovered at 0213. Upon recovery it was noticed that
the Seatech Fluorometer had not been removed prior to the deep cast.
The data seemed OK and M. Valenciano inspected the inside of the
electronics housing and found it to be dry. We performed an additional
test cast to 1000 m at 0923 and found the fluorometer uptrace data to
be dubious. The unit will be sent back to WetLabs for inspection and
The sediment trap array was recovered at around 0645. The array had
drifted to the west. After the sediment trap array had been recovered
we steamed back to Station ALOHA.
Upon arrival at Station ALOHA, the ATE sampler was deployed at 0923.
The PRR and TSRB were deployed at 1202 and AC-9/FRRf casts were
performed at 0306, 1235, and 1352.
We conducted a 2500 m cast at Station Kaena however we abbreviated the
Nisken bottle sampling scheme due to a medical emergency. This cast
was recovered at 2152 at which time we began transit to Snug Harbor.
November 6, 2002
Arrived at Snug Harbor at around 0330. A full offload took place later
Sub component programs:
Bob Bidigare HPLC pigments/UH
Mike Landry zooplankton dynamics/UH
John Dore CO2 dynamics/UH
Charles Keeling CO2 dynamics and intercalibration/SIO
Mark Abbott/Ricardo Letelier optical measurements/OSU
Penny Chisholm/Erik Zinser Prochlorococcus ecotype dynamics/MIT